# Simulations of Ellipsoidal Primordial Black Hole Formation

### v1

@@ 1 Introduction
Non-spherical PBH formation during a radiation-dominated Universe was studied in [18] through $3+1$ numerical simulations, considering a case with spheroidal ($p=\pm e$) initial curvature fluctuation. The results showed that, for a typical fluctuation amplitude and a Gaussian spatial profile, the threshold for non-spherical collapse is not significantly modified compared with the spherical case. In this work, we aim to consider a different curvature profile sourced by a monochromatic power spectrum following peak theory and determine the threshold values for $e$ and $p$. We further explore how these thresholds vary with different fluctuation amplitudes Clarifying the PBH formation condition including the parameters $e$ and $p$ is crucial for statistically estimating the impact of non-spherical effects on the PBH mass function. This can also be a starting point for the generalization of analytical threshold estimations for non-spherical collapse, taking into account the values of $e$ and $p$ as the parameters characterizing the initial profile, as done in [28, 29] in the case of spherical symmetry.
@@
Non-spherical effects may be relevant for some physically well-motivated scenarios where soft equations of state are considered and predictions of the mass function of PBHs may be contrasted with the gravitational wave spectrum of induced gravitational waves (see [39] for a review) and/or with merger rates of PBH binaries [40]. The softening can be due to a still unknown phase transition or to the standard thermal history of the Universe [41] when particles become non-relativistic. For instance, the softening of the equation of state due to the QCD crossover about $\sim 10^{-5}{\rm s}$ after the Big Bang [42, 43, 44] might cause a significant number of PBHs in the solar mass range [45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. Another possibility is from crossovers beyond standard model theories that may induce a significant softening of the equation of state [55, 56]. In addition, several scenarios with $w\neq 1/3$ have also been considered in the context of the PTA signal [57, 58] and its counterpart with PBHs [59, 60, 61, 62, 63].
@@ 2 Deviation from sphericity of the initial curvature perturbation in peak theory
$\frac{\bar{\zeta}}{\sigma_{0}}=\frac{\nu}{1-\gamma^{2}}\left(\psi+R^{2}_{s}\frac{\nabla^{2}\psi}{3}\right)-\frac{\xi/\gamma}{(1-\gamma^{2})}\left(\gamma^{2}\psi+\frac{R^{2}_{s}\nabla^{2}\psi}{3}\right)+\frac{5}{2}R^{2}_{s}\left(\frac{\xi}{\gamma}\right)\left(\frac{\psi^{\prime}}{r}-\frac{\nabla^{2}\psi}{3}\right)^{2}A(e,p),$ (2.15)
@@ 3 Numerical method and procedure
where $\alpha$ is the lapse function, $\beta^{i}$ is the shift-vector, $\tilde{\gamma}_{ij}$ is the conformal metric whose determinant is given by that of the flat reference metric, $\tilde{\psi}$ is the spatial conformal factor, which at super-horizon scales is $\psi=a^{1/2}\Psi=a^{1/2}\exp(-\zeta/2)$ with $a$ being the scale factor of the FLRW background. All these variables are functions of $t$ and $\vec{x}$. The time evolution of these variables at super-horizon scales has been analytically investigated in Ref. [67]. The energy momentum-tensor is given by the following perfect fluid form:
@@ 4.1 Dynamical evolution of the non-spherical gravitational collapse
where $\tilde{p}_{\rm c}^{\pm}(e)$ is defined such that the critical line is given by $p=\tilde{p}_{\rm c}(e)=\tilde{p}_{\rm c}^{\pm}(e)$ for $p\gtrless 0$, and $e_{0}$, $p^{\pm}_{0}$ and the exponent $n^{\pm}$ are the parameters characterising the superellipse. Making a non-linear fit of Eq. (4.1) with our numerical results, we obtain $n^{+}\approx 2.48$, $n^{-}\approx 1.74$ for $w=1/3$ and $n^{+}\approx 2.52$, $n^{-}\approx 1.67$ for $w=1/10$. The behaviour of $\tilde{p}_{\rm c}(e)$ is not symmetric for the reflection along $e$ axis, that is, $\tilde{p}_{\rm c}^{+}(e)\neq-\tilde{p}_{\rm c}^{-}(e)$. This is expected since the set of eigenvalues $\lambda_{l}$ is not the same when considering $p\rightarrow-p$ with a fixed $e$. For the cases $e=0$, we found $\tilde{p}_{\rm c}^{+}(0)=p_{0}^{+}<|\tilde{p}_{\rm c}^{-}(0)|=|p_{0}^{-}|$ indicating that a slightly larger deviation from sphericity is required for the fluctuations to collapse for prolate cases ($p<0$) compared with oblate cases ($p>0$). Note that the value of $\tilde{p}_{\rm c}(e)$ does not change much in the region where the probability given by Eq. (2.11) is non-zero. This supports the idea that we can ignore the $p$ dependence in the threshold estimation when we estimate the PBH mass function.
@@ 4.3 Critical ellipticity as a function of the initial amplitude $\mu$
Figure 12: “Phase diagram” in the $(e,p)$ plane for the typical initial amplitude $\mu=\mu_{t}$. The green and red dots denote configurations of $e$ and $p$ that avoid and form black holes, respectively. The magenta and orange lines represent the numerical fitting of Eq.(4.1) to the numerical threshold values (infer as the middle point between the green and red points) for the case $w=1/10$ and $w=1/3$, respectively. The coloured region corresponds to the probability distribution $\mathcal{P}_{e,p}$ (Eq.(2.11)) fixing $\nu_{c}=8$.
@@
On the other hand, if we focus on the critical scaling regime, in which the amplitude is very close to $\mu_{\rm c,sp}$, the number of configurations collapsing into black holes can be significantly reduced. For instance, with a shift of $0.1\%$ in the threshold of $\mu_{\rm c,e}$ compared with $\mu_{\rm c,sp}$, most of the configurations are prevented from collapsing into black holes although it slightly depends on $w$ and $\nu$. Therefore the non-spherical effects can be highly significant in the critical regime, preventing a large fraction of configurations from collapsing into black holes. Nevertheless, we report in the accompanying letter [82] that, in terms of the PBH mass function, the effects of the ellipticity may not have a significant impact even on the power-law small mass tail originates from the critical behavior.
@@ 5 Conclusions
In this work, we employed $3+1$ relativistic numerical simulations to study the collapse of super-horizon curvature fluctuations with an ellipsoidal geometry, characterized by ellipticity ($e$) and prolateness ($p$), in line with peak theory [26] assuming a monochromatic power spectrum. When analyzing the dynamics for the two cases of the equation of state with $w=1/10$ and $1/3$, we observe a characteristic behaviour of the oscillating ellipsoidal shape between oblate and prolate configurations. This oscillation persists until very late times, when the shape becomes nearly spherical, just before the formation of the apparent horizon. Although this implies that the assumption of an exactly spherical shape is not valid until the final stages of collapse, our results indicate that, for the cases tested, the non-sphericity decays over time, in agreement with [80, 22], with relatively small non-sphericities in the initial data.
@@ Appendix B High peaks for PBH formation assuming spherical symmetry
Figure 15: Top panels: PBH mass functions for several values of $w$ with the monochromatic power spectrum as functions of $M/M_{k_{p}}$ (left) and $\mu-\mu_{c}$ (right). We fix $k_{p}=10^{13.5}\textrm{Mpc}^{-1}$ (asteroid mass range) with $g_{*}=106.75$ in both cases. The dashed vertical lines in the left panel denote the typical values $\mu_{t}$ for $w=1/10$ and $w=1/3$ used in the simulations. Middle panels: Critical values of $\nu_{c}$ (left) and the amplitude of the power spectrum $\mathcal{A}_{\zeta}$ (right) as functions of $w$. The solid and dashed lines correspond to $k_{p}=10^{13.5}$ (asteroid mass range) and $k_{p}=10^{7}$ (solar mass range), respectively, with $g_{*}=10.75$. Bottom panel: Value of $f_{\rm PBH}^{\rm tot}$ as a function of $\mu_{c}$ for $k_{p}=10^{13.5}\textrm{Mpc}^{-1}$.

### v2

@@ 1 Introduction
Non-spherical PBH formation during a radiation-dominated Universe was studied in [18] through $3+1$ numerical simulations, considering a case with spheroidal ($p=\pm e$) initial curvature fluctuation. The results showed that, for a typical fluctuation amplitude and a Gaussian spatial profile, the threshold for non-spherical collapse is not significantly modified compared with the spherical case. In this work, we aim to consider a different curvature profile sourced by a monochromatic power spectrum following peak theory and determine the threshold values for $e$ and $p$. We further explore how these thresholds vary with different fluctuation amplitudes. Clarifying the PBH formation condition including the parameters $e$ and $p$ is crucial for statistically estimating the impact of non-spherical effects on the PBH mass function. This can also be a starting point for the generalization of analytical threshold estimations for non-spherical collapse, taking into account the values of $e$ and $p$ as the parameters characterizing the initial profile, as done in [28, 29] in the case of spherical symmetry.
@@
Non-spherical effects may be relevant for some physically well-motivated scenarios where soft equations of state are considered and predictions of the mass function of PBHs may be contrasted with the gravitational wave spectrum of induced gravitational waves (see [39] for a review) and/or with merger rates of PBH binaries [40, 41]. The softening can be due to a still unknown phase transition or to the standard thermal history of the Universe [42] when particles become non-relativistic. For instance, the softening of the equation of state due to the QCD crossover about $\sim 10^{-5}{\rm s}$ after the Big Bang [43, 44, 45] might cause a significant number of PBHs in the solar mass range [46, 47, 48, 49, 50, 51, 52, 53, 54, 55]. Another possibility is from crossovers beyond standard model theories that may induce a significant softening of the equation of state [56, 57]. In addition, several scenarios with $w\neq 1/3$ have also been considered in the context of the PTA signal [58, 59] and its counterpart with PBHs [60, 61, 62, 63, 64, 65, 66].
@@ 2 Deviation from sphericity of the initial curvature perturbation in peak theory
$\frac{\bar{\zeta}}{\sigma_{0}}=\frac{\nu}{1-\gamma^{2}}\left(\psi+R^{2}_{s}\frac{\nabla^{2}\psi}{3}\right)-\frac{\xi/\gamma}{(1-\gamma^{2})}\left(\gamma^{2}\psi+\frac{R^{2}_{s}\nabla^{2}\psi}{3}\right)+\frac{5}{2}R^{2}_{s}\left(\frac{\xi}{\gamma}\right)\left(\frac{\psi^{\prime}}{r}-\frac{\nabla^{2}\psi}{3}\right)A(e,p),$ (2.15)
@@ 3 Numerical method and procedure
where $\alpha$ is the lapse function, $\beta^{i}$ is the shift-vector, $\tilde{\gamma}_{ij}$ is the conformal metric whose determinant is given by that of the flat reference metric, $\tilde{\psi}$ is the spatial conformal factor, which at super-horizon scales is $\tilde{\psi}=a^{1/2}\Psi=a^{1/2}\exp(-\zeta/2)$ with $a$ being the scale factor of the FLRW background. All these variables are functions of $t$ and $\vec{x}$. The time evolution of these variables at super-horizon scales has been analytically investigated in Ref. [70]. The energy momentum-tensor is given by the following perfect fluid form:
@@ 4.1 Dynamical evolution of the non-spherical gravitational collapse
where $\tilde{p}_{\rm c}^{\pm}(e)$ is defined such that the critical line is given by $p=\tilde{p}_{\rm c}(e)=\tilde{p}_{\rm c}^{\pm}(e)$ for $p\gtrless 0$, and $e_{0}$, $p^{\pm}_{0}$ and the exponent $n^{\pm}$ are the parameters characterising the superellipse. Making a non-linear fit of Eq. (4.1) with our numerical results, we obtain $n^{+}\approx 2.44\pm 0.03$, $n^{-}\approx 1.70\pm 0.02$ for $w=1/3$ and $n^{+}\approx 2.53\pm 0.10$, $n^{-}\approx 1.70\pm 0.05$ for $w=1/10$. The behaviour of $\tilde{p}_{\rm c}(e)$ is not symmetric for the reflection along $e$ axis, that is, $\tilde{p}_{\rm c}^{+}(e)\neq-\tilde{p}_{\rm c}^{-}(e)$. This is expected since the set of eigenvalues $\lambda_{l}$ is not the same when considering $p\rightarrow-p$ with a fixed $e$. For the cases $e=0$, we found $\tilde{p}_{\rm c}^{+}(0)=p_{0}^{+}<|\tilde{p}_{\rm c}^{-}(0)|=|p_{0}^{-}|$ indicating that a slightly larger deviation from sphericity is required for the fluctuations to collapse for prolate cases ($p<0$) compared with oblate cases ($p>0$). Note that the value of $\tilde{p}_{\rm c}(e)$ does not change much in the region where the probability given by Eq. (2.11) is non-zero. This supports the idea that we can ignore the $p$ dependence in the threshold estimation when we estimate the PBH mass function.
@@ 4.3 Critical ellipticity as a function of the initial amplitude $\mu$
Figure 12: “Phase diagram” in the $(e,p)$ plane for the typical initial amplitude $\mu=\mu_{t}$. The red and green dots denote configurations of $e$ and $p$ that avoid and form black holes, respectively. The magenta and orange lines represent the numerical fitting of Eq.(4.1) to the numerical threshold values (infer as the middle point between the green and red points) for the case $w=1/10$ and $w=1/3$, respectively. The coloured region corresponds to the probability distribution $\mathcal{P}_{e,p}$ (Eq.(2.11)) fixing $\nu_{c}=8$.
@@
On the other hand, if we focus on the critical scaling regime, in which the amplitude is very close to $\mu_{\rm c,sp}$, the number of configurations collapsing into black holes can be significantly reduced. For instance, for $(\mu-\mu_{\rm c,sp})/\mu_{\rm c,sp}=10^{-3}$, most of the configurations are prevented from collapsing into black holes although it slightly depends on $w$ and $\nu$. Therefore the non-spherical effects can be highly significant in the critical regime, preventing a large fraction of configurations from collapsing into black holes. Nevertheless, we report in the accompanying letter [85] that, in terms of the PBH mass function, the effects of the ellipticity may not have a significant impact even on the power-law small mass tail originates from the critical behavior.
@@ 5 Conclusions
In this work, we employed $3+1$ relativistic numerical simulations to study the collapse of super-horizon curvature fluctuations with an ellipsoidal geometry, characterized by ellipticity ($e$) and prolateness ($p$), in line with peak theory [26] assuming a monochromatic power spectrum. When analyzing the dynamics for the two cases of the equation of state with $w=1/10$ and $1/3$, we observe a characteristic behaviour of the oscillating ellipsoidal shape between oblate and prolate configurations. This oscillation persists until very late times, when the shape becomes nearly spherical, just before the formation of the apparent horizon. Although this implies that the assumption of an exactly spherical shape is not valid until the final stages of collapse, our results indicate that, for the cases tested, the non-sphericity decays over time, in agreement with [83, 22], with relatively small non-sphericities in the initial data. On the other hand, non-spherical collapse will be accompanied by the emission of gravitational waves [90]; however, a thorough investigation of this phenomenon is left for future research.
@@ Appendix B High peaks for PBH formation assuming spherical symmetry
Figure 15: Top panels: PBH mass functions for several values of $w$ with the monochromatic power spectrum as functions of $M/M_{k_{p}}$ (left) and $\mu-\mu_{c}$ (right) with $M_{k_{p}}=(g_{\star}/106.75)^{-1/6}\cdot 10^{20}\left(k_{p}/(1.56\cdot 10^{13}\textrm{Mpc}^{-1})\right)^{-3(1+w)/(1+3w)}M_{\odot}$. We fix $k_{p}=10^{13.5}\textrm{Mpc}^{-1}$ (asteroid mass range) with $g_{*}=106.75$ in both cases. The dashed vertical lines in the left panel denote the typical values $\mu_{t}$ for $w=1/10$ and $w=1/3$ used in the simulations. Middle panels: Critical values of $\nu_{c}$ (left) and the amplitude of the power spectrum $\mathcal{A}_{\zeta}$ (right) as functions of $w$. The solid and dashed lines correspond to $k_{p}=10^{13.5}$ (asteroid mass range) and $k_{p}=10^{7}$ (solar mass range), respectively, with $g_{*}=10.75$. Bottom panel: Value of $f_{\rm PBH}^{\rm tot}$ as a function of $\mu_{c}$ for $k_{p}=10^{13.5}\textrm{Mpc}^{-1}$.